OpenStudy (anonymous):

9c(c-11)+10(5c-3)=3c(c+5)+c(6c-3)-30 I got -49c=12c but I know that it's not right please help

OpenStudy (nikato):

I don't think so. There's either no answer or c=0

OpenStudy (anonymous):

9c(c-11)+10(5c-3)=3c(c+5)+c(6c-3)-30 Well first you should distribute all the brackets in the equation by multiplying in whats inside. 9c-99c+50c-30 = 3c+15c+6c-3c-30 THEN, we just simplify, by combining all like terms! -90c+50c-30=28c+6c-3c-30 -40c-30=31c-30 Subtract 31 from both sides. -40c-30-31c=31c-30-31c Add positive 30 to both sides. -40c-30-31c+30=-30+30 -71c=0 Divide -71. C = 0

OpenStudy (anonymous):

OMG thank you so much!